OpenStudy (anonymous):

Solve using substitution or elimination method 3x - 2y = 15 6x + 5y = 11

OpenStudy (anonymous):

@pooja195 @sleepyjess

pooja195 (pooja195):

elimination might be easier here

OpenStudy (anonymous):

that is what i think

pooja195 (pooja195):

ok multiply the top equation by -2 to make the equation opposite

pooja195 (pooja195):

\[-2(3x - 2y = 15)\]

OpenStudy (anonymous):

see here is my question Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 ------- 3x - 2y = 15 6x + 5y = 11 ------- 4x - 3y = 9 -8x + 6y = -18 ------- 5x - 3y = 16 -2x + 3y = -7 i know A isn't it, but i am trying to solve for B using the elmination method. last people tried but gave some weird hard explination not using any of those two methods

OpenStudy (anonymous):

ok that is what I did at first, so that means 3x - 2y = 15 -> 6x + 4y = -30 right

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@iambatman help???

pooja195 (pooja195):

nope and sorry for the late response this site it laggy >_<

OpenStudy (anonymous):

wait would the 4y be -4y

OpenStudy (anonymous):

can we make a twiddla or something to work on?

pooja195 (pooja195):

pm me the link ill help u there

OpenStudy (anonymous):

would be a whole lot easier than here right now

pooja195 (pooja195):

-6x+4y=-30